OpenStudy (idealist10):

How to integrate -2x^3*e^x^2 dx?

OpenStudy (idealist10):

\[\int\limits_{}^{}-2x^3e ^{x^2} dx\]

OpenStudy (idealist10):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

e^(x^2) has no elementary antiderivative so we cannot integrate e^(x^2) and get a known function type (eg: radical, trig, etc) but -2x^3 can be integrated just fine

jimthompson5910 (jim_thompson5910):

use integration by parts let u = e^(x^2) so that means du = 2x*e^(x^2)dx and let dv = -2x^3 integrating both sides gives v = (-x^4)/2

OpenStudy (anonymous):

just in case you didn't know (jim didn't explicitly say it), the derivative of e^u=u' * e^u

OpenStudy (zarkon):

\[x^3e^{x^2}=x^2xe^{x^2}\]

OpenStudy (agent0smith):

\[\Large \int\limits\limits_{}^{}-2x^3e ^{x^2} dx = - \int\limits x^2*2x e^{x^2} dx\]

jimthompson5910 (jim_thompson5910):

doing it by parts is messy, but we can use this trick to make things a bit simpler -2x^3*e^(x^2) -2x*x^2*e^(x^2) then we can use u-sub u = x^2 du = 2xdx so we'll end up with -u*e^(u)du now integration by parts shouldn't be so bad

OpenStudy (idealist10):

Not so bad by using this trick. Thank you!

jimthompson5910 (jim_thompson5910):

no problem